OpenStudy (anonymous):

Given the equation P2 = A3, what is the orbital period, in years, for the planet Saturn? (Saturn is located 9.5 AU from the sun.) 5 years 29 years 229 years 500 years

OpenStudy (anonymous):

If \(p^2=a^3\) then \(p=a^{1.5}\). Since you know a it is just plugging in.